Gina is 9 years younger than her sister, Bridgette. Two years ago, Bridgette was twice as old as Gina was. How old is Gina now?
Answer by checkley77(12844)

GINA NOW=X
BRIDGETTE=X+9
GINA 2 YEARS AGO=X-2
BRIDGETTE 2 YEARS AGO=(X+9-2)
(X+9-2)=2(X-2)
X+7=2X-4
X-2X=-4-7
-X=-11
X=11 ANS. FOR GINA'S AGE NOW.
11+9=20 ANS. FOR BRIDGETTE'S AGE NOW.
PROOF:
20-2=2(11-2)
18=2*9
18=18